# MMLU-Pro+: Evaluating Higher-Order Reasoning and Shortcut Learning in LLMs

### v1

@@ Abstract
Existing benchmarks for large language models (LLMs) increasingly struggle to differentiate between top-performing models, underscoring the need for more challenging evaluation frameworks. We introduce MMLU-Pro+, an enhanced benchmark building upon MMLU-Pro to assess shortcut learning and higher-order reasoning in LLMs. By incorporating questions with multiple correct answers across diverse domains, MMLU-Pro+ tests LLMs’ ability to engage in complex reasoning and resist simplistic problem-solving strategies. Our results show that MMLU-Pro+ maintains MMLU-Pro’s difficulty while providing a more rigorous test of model discrimination, particularly in multi-correct answer scenarios. We introduce novel metrics like shortcut selection ratio and correct pair identification ratio, offering deeper insights into model behavior and anchoring bias. Evaluations of five state-of-the-art LLMs reveal significant performance gaps, highlighting variations in reasoning abilities and bias susceptibility. We release the dataset and evaluation codes at https://github.com/asgsaeid/mmlu-pro-plus.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have led to remarkable improvements in various natural language processing tasks [7, 15]. State-of-the-art models such as GPT-4 [1], Claude-3.5 Sonnet [2], Gemini [17], and Llama [20] have demonstrated impressive capabilities across a wide range of applications. However, as these models continue to evolve, existing benchmarks are struggling to keep pace, often reaching performance saturation and failing to effectively differentiate between model capabilities [10].
@@ 3 Experiments
We evaluated several state-of-the-art LLMs, including two open-source models (LLaMA-3.1 405B Instruct [20] and Qwen 2 72B Instruct [9]) and three closed-source models (GPT-4o [1], Claude Sonnet 3.5 [2], and Gemini-1.5 Pro [17]), using the MMLU-Pro+ benchmark. Models were assessed not only on accuracy but also on their ability to recognize and correctly select the "Both X and Y are correct" option, thereby demonstrating higher-order reasoning and shortcut leanring.
@@ 3.1 Accuracy Analysis on MMLU-Pro+
Sonnet-3.5 demonstrates superior performance across the majority of categories, achieving the highest average accuracy. The consistent outperformance suggests a more robust capability in handling the increased complexity introduced by MMLU-Pro+. The universal decrease in accuracy from MMLU-Pro to MMLU-Pro+ validates the increased challenge of our new benchmark. Notably, Sonnet-3.5 exhibits the smallest average performance drop, indicating enhanced resilience to the modified question format and the introduction of multiple correct answers.
@@
The engineering category reveals an interesting divergence, with Gemini-1.5-Pro and Sonnet-3.5 showing markedly smaller performance drops compared to other models. Conversely, the law category presents an anomaly, with minimal performance drops and even slight improvements for some models, suggesting that the MMLU-Pro+ modifications may have had less impact on this domain. Despite high scores in many categories, GPT-4o exhibits the largest average performance drop, indicating a potential sensitivity to the structural changes in MMLU-Pro+. This suggests that high performance on standard benchmarks may not necessarily translate to robust reasoning capabilities in more complex scenarios.
@@
The performance gap between different models on MMLU-Pro, particularly the superior performance of models like Claude-3.5-Sonnet compared to GPT-4o (which was used in dataset creation), further validates our dataset construction methodology. Despite GPT-4o’s involvement in generating True Positive pairs, it does not exhibit an advantage in the question-answering task. This discrepancy in performance across models suggests that MMLU-Pro successfully challenges the LLMs, requiring genuine reasoning capabilities rather than pattern matching or exploitation of dataset artifacts.
| Category | Qwen-2-72B-Ins | Gemini-1.5-Pro | GPT-4o | Llama-405B-Ins | Sonnet-3.5 |
| biology | 72.3-9.7 | 73.8-9.9 | 77.8-10.4 | 79.2-6.0 | 79.6-7.8 |
| business | 45.8-22.2 | 55.5-16.6 | 53.1-26.5 | 59.4-17.6 | 67.4-12.2 |
| chemistry | 42.5-16.3 | 52.4-17.7 | 49.9-25.7 | 57.8-15.1 | 64.3-12.3 |
| computer science | 49.5-18.5 | 54.4-14.5 | 56.6-23.6 | 60.5-13.9 | 65.9-14.9 |
| economics | 63.3-13.3 | 62.8-13.9 | 71.4-11.1 | 70.6-9.8 | 73.9-8.5 |
| engineering | 37.8-9.8 | 40.7-3.9 | 37.3-18.5 | 46.3-13.5 | 55.3-4.2 |
| health | 58.6-8.4 | 63.0-4.4 | 67.2-8.1 | 65.8-6.5 | 70.7-6.1 |
| history | 60.4-6.0 | 58.8-7.3 | 70.1-2.4 | 60.9-6.6 | 71.9-1.3 |
| law | 43.6-0.7 | 47.8–0.5 | 51.4-3.3 | 55.4–1.2 | 56.9-7.3 |
| math | 47.2-23.4 | 53.4-8.8 | 52.1-25.9 | 61.5-15.7 | 66.5-9.8 |
| other | 60.1-6.0 | 59.8-10.3 | 68.1-9.8 | 67.4-5.7 | 71.4-6.7 |
| philosophy | 51.1-8.2 | 51.7-11.3 | 64.8-6.8 | 61.4-4.8 | 66.3-8.2 |
| physics | 43.2-18.2 | 54.3-14.9 | 53.5-21.6 | 58.0-14.2 | 63.2-13.4 |
| psychology | 69.8-6.4 | 66.8-9.7 | 75.4-5.9 | 72.5-4.8 | 76.9-5.6 |
| Average | 53.2-11.9 | 56.8-10.2 | 60.6-14.3 | 62.6-9.5 | 67.9-8.5 |
@@ 3.2 Analysis of Anchoring Bias and Shortcut Learning in MMLU-Pro+
The persistence in selecting previously incorrect options (high ) is especially noteworthy, as it indicates potential limitations in these models’ ability to reassess and engage in higher-order reasoning when presented with new, valid alternatives. Similarly, a high suggests a failure to recognize newly introduced correct options. Conversely, Gemini 1.5 Pro, Sonnet-3.5, and Llama 1.3 405B demonstrate lower shortcut selection ratios, suggesting a greater capacity for adapting their reasoning in light of new information. These findings highlight the challenges language models face in fully leveraging the additional correct options introduced in MMLU-Pro+, and underscore the importance of developing benchmarks that can effectively evaluate and promote higher-order reasoning.
@@ 5 Discussion
Model Differentiation. MMLU-Pro+ revealed substantial differences in model performance, with Sonnet-3.5 consistently outperforming other models across most categories.
Anchoring Bias and Shortcut Learning. The shortcut selection ratio analysis exposed varying degrees of anchoring bias across models, highlighting the challenge LLMs face in adapting their reasoning when presented with new, valid alternatives.
Higher-Order Reasoning. The correct pair identification ratio provided insights into models’ abilities to distinguish genuinely correct answer pairs from misleading options, with significant variations observed across different models.
These findings underscore the importance of sophisticated evaluation metrics in assessing advanced language models, particularly in scenarios requiring discernment between subtly different correct and incorrect information combinations.

### v2

@@ Abstract
Existing benchmarks for large language models (LLMs) increasingly struggle to differentiate between top-performing models, underscoring the need for more challenging evaluation frameworks. We introduce MMLU-Pro+, an enhanced benchmark building upon MMLU-Pro to assess shortcut learning and higher-order reasoning in LLMs. By incorporating questions with multiple correct answers across diverse domains, MMLU-Pro+ tests LLMs’ ability to engage in complex reasoning and resist simplistic problem-solving strategies. Our results show that MMLU-Pro+ maintains MMLU-Pro’s difficulty while providing a more rigorous test of model discrimination, particularly in multi-correct answer scenarios. We introduce novel metrics like shortcut selection ratio and correct pair identification ratio, offering deeper insights into model behavior and anchoring bias. Evaluations of six state-of-the-art LLMs reveal significant performance gaps, highlighting variations in reasoning abilities and bias susceptibility. We release the dataset and evaluation codes at https://github.com/asgsaeid/mmlu-pro-plus.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have led to remarkable improvements in various natural language processing tasks [7, 15]. State-of-the-art models such as GPT-4o, O1-preview, [1], Claude-3.5 Sonnet [2], Gemini [17], and Llama [20] have demonstrated impressive capabilities across a wide range of applications. However, as these models continue to evolve, existing benchmarks are struggling to keep pace, often reaching performance saturation and failing to effectively differentiate between model capabilities [10].
@@ 3 Experiments
We evaluated several state-of-the-art LLMs, including two open-source models (LLaMA-3.1-405B-Instruct [20] and Qwen-2-72B-Instruct [9]) and four closed-source models (OpenAI’s GPT-4o [1] and O1-preview, Claude-Sonnet-3.5 [2], and Gemini-1.5-Pro [17]), using the MMLU-Pro+ benchmark. Models were assessed not only on accuracy but also on their ability to recognize and correctly select the "Both X and Y are correct" option and avoid partially correct options, thereby demonstrating higher-order reasoning and shortcut leanring.
@@ 3.1 Accuracy Analysis on MMLU-Pro+
O1-preview demonstrates superior performance across the majority of categories, achieving the highest average accuracy. The consistent outperformance suggests a more robust capability in handling the increased complexity introduced by MMLU-Pro+. The universal decrease in accuracy from MMLU-Pro to MMLU-Pro+ validates the increased challenge of our new benchmark. Notably, O1-preview exhibits the smallest average performance drop, indicating enhanced resilience to the modified question format and the introduction of multiple correct answers.
@@
The engineering category reveals an interesting divergence, with Gemini-1.5-Pro and Sonnet-3.5 showing markedly smaller performance drops compared to other models. Conversely, the law category presents an anomaly, with minimal performance drops and even slight improvements for some models, suggesting that the MMLU-Pro+ modifications may have had less impact on this domain. GPT-4o exhibits the largest average performance drop, indicating a potential sensitivity to the structural changes in MMLU-Pro+. This suggests that high performance on standard benchmarks may not necessarily translate to robust reasoning capabilities in more complex scenarios.
@@
The performance gap between different models on MMLU-Pro, particularly the superior performance of models like O1-preview and Claude-3.5-Sonnet compared to GPT-4o (which was used in dataset creation), further validates our dataset construction methodology. Despite GPT-4o’s involvement in generating True Positive pairs, it does not exhibit an advantage in the question-answering task. This discrepancy in performance across models suggests that MMLU-Pro successfully challenges the LLMs, requiring genuine reasoning capabilities rather than pattern matching or exploitation of dataset artifacts.
| Category | Qwen2-72B-Ins | Gemini-1.5-Pro | GPT-4o | Llama-405B-Ins | Sonnet-3.5 | O1-preview |
| biology | 72.3-9.7 | 73.8-9.9 | 77.8-10.4 | 79.2-6.0 | 79.6-7.8 | 79.3-9.9 |
| business | 45.8-22.2 | 55.5-16.6 | 53.1-26.5 | 59.4-17.6 | 67.4-12.2 | 64.3-23.7 |
| chemistry | 42.5-16.3 | 52.4-17.7 | 49.9-25.7 | 57.8-15.1 | 64.3-12.3 | 68.5-17.2 |
| computer science | 49.5-18.5 | 54.4-14.5 | 56.6-23.6 | 60.5-13.9 | 65.9-14.9 | 67.9–67.9 |
| economics | 63.3-13.3 | 62.8-13.9 | 71.4-11.1 | 70.6-9.8 | 73.9-8.5 | 76.1-9.6 |
| engineering | 37.8-9.8 | 40.7-3.9 | 37.3-18.5 | 46.3-13.5 | 55.3-4.2 | 53.0-15.6 |
| health | 58.6-8.4 | 63.0-4.4 | 67.2-8.1 | 65.8-6.5 | 70.7-6.1 | 71.4-7.7 |
| history | 60.4-6.0 | 58.8-7.3 | 70.1-2.4 | 60.9-6.6 | 71.9-1.3 | 65.1-9.4 |
| law | 43.6-0.7 | 47.8–0.5 | 51.4-3.3 | 55.4–1.2 | 56.9-7.3 | 57.0-11.5 |
| math | 47.2-23.4 | 53.4-8.8 | 52.1-25.9 | 61.5-15.7 | 66.5-9.8 | 71.4-18.8 |
| other | 60.1-6.0 | 59.8-10.3 | 68.1-9.8 | 67.4-5.7 | 71.4-6.7 | 72.5-8.5 |
| philosophy | 51.1-8.2 | 51.7-11.3 | 64.8-6.8 | 61.4-4.8 | 66.3-8.2 | 67.5-12.0 |
| physics | 43.2-18.2 | 54.3-14.9 | 53.5-21.6 | 58.0-14.2 | 63.2-13.4 | 69.4-17.7 |
| psychology | 69.8-6.4 | 66.8-9.7 | 75.4-5.9 | 72.5-4.8 | 76.9-5.6 | 73.1-11.8 |
| Average | 53.2-11.9 | 56.8-10.2 | 60.6-14.3 | 62.6-9.5 | 67.9-8.5 | 68.3-7.5 |
@@ 3.2 Analysis of Anchoring Bias and Shortcut Learning in MMLU-Pro+
The persistence in selecting previously incorrect options (high ) is especially noteworthy, as it indicates potential limitations in these models’ ability to reassess and engage in higher-order reasoning when presented with new, valid alternatives. Similarly, a high suggests a failure to recognize newly introduced correct options. Conversely, Gemini 1.5 Pro, Sonnet-3.5, and Llama 1.3 405B demonstrate lower shortcut selection ratios, suggesting a greater capacity for adapting their reasoning in light of new information. Interestingly, O1-preview has improved compared to GPT-4o, but still lags behind Gemini-1.5-Pro and Sonnet-3.5. These findings highlight the challenges language models face in fully leveraging the additional correct options introduced in MMLU-Pro+, and underscore the importance of developing benchmarks that can effectively evaluate and promote higher-order reasoning.
@@ 5 Discussion
Model Differentiation. MMLU-Pro+ revealed substantial differences in model performance. While O1-preview outperformed other models across most categories in general assessments, its dominance diminishes when anchoring bias and higher-order reasoning are considered.
Anchoring Bias and Shortcut Learning. The shortcut selection ratio analysis exposed varying degrees of anchoring bias across models, highlighting the challenge LLMs face in adapting their reasoning when presented with new, valid alternatives. Notably, Gemini-1.5-Pro and Sonnet-3.5 demonstrated less reliance on shortcuts.
Higher-Order Reasoning. The correct pair identification ratio provided insights into models’ abilities to distinguish genuinely correct answer pairs from misleading options, with significant variations observed across different models. Sonnet-3.5 significantly outperformed others in this aspect.
These findings underscore the importance of new evaluation metrics in assessing advanced language models, particularly in scenarios requiring discernment between subtly different correct and incorrect information combinations.